OpenStudy (anonymous):

How would you use the distributive property to solve: 3p - 6q = 3(__?__)

OpenStudy (airyana1114):

Okay, so do you know what the distributive property is?

OpenStudy (anonymous):

Of course :)

OpenStudy (anonymous):

a(b + c) = ab + ac

OpenStudy (airyana1114):

Okay, so then we're basically working backwards. How do you think they got 3p from a 3?

OpenStudy (anonymous):

Ooooh! I see! It's p - 2q, correct?

OpenStudy (airyana1114):

Yes :) Good job!

OpenStudy (anonymous):

But this is a proof, so I have to show each step. I just did that by reasoning... What would the steps be? I'll give you a medal already tho. :) Thank you!

OpenStudy (airyana1114):

Haha, okay thank you so I'm gonna take a few minutes and show you how, step by step. Is that okay?

OpenStudy (anonymous):

Yup! :D

OpenStudy (airyana1114):

Step 1: Divide 3p by 3. This will give you 3p - 6q = 3(p?__) Step 2: DIvide -6q by 3. This will give you 3p - 6q = 3(p - 2q)

OpenStudy (airyana1114):

Does that clear it up for you? So if you have this problem again, you just divide by the number outside of the parentheses since we're doing the opposite of distributing. (Dividing.)

OpenStudy (anonymous):

Yep! Thanks so much! :)

OpenStudy (airyana1114):

You're welcome :)